a bin contains 7 red cards,9green cards 3 yellow cards and 6 blue cards .find the probability of selecting 3 green cards one at a time without replacement
evaluate this expression 17p4

Answer by stanbon(75887)

a bin contains 7 red cards,9green cards 3 yellow cards and 6 blue cards .find the probability of selecting 3 green cards one at a time without replacement
---
P(3 green) = 9C3/25C3 = 0.0365
--------------------------------------
evaluate this expression 17p4 = (17*16*15*14)/(1*2*3*4) = 2380
==========================
Cheers,
Stan H.